1. A firm’s MC//MR at 30 units of output and a price of $12. ATC at 30 units of output is $15.00, AVC are $13, and consumers are demanding 30 units at a price of $20. What is the economic outcome? What type of profit/loss does this represent to the firm? What is the total profit/loss? Show your calculations.
Answer #1

Answer:

1. Here firm is producing 30 units when MR = MC, it means firm is maximizing its profit. Thus output is profit maximizing out put

2. Here firm is charging price higher than marginal revenue means firm is in imperfect market

3. Firm is earning economic profit because firm is getting more profit than break even

4. Economic profit = ( P - ATC ) * Q = ( 20 - 15 ) * 30 = 150